OpenStudy (anonymous):

rewrite as a single log 2ln(x+1) + 1/3lnx - ln(cosx)

OpenStudy (anonymous):

\[=\ln(x+1)^2+\ln \sqrt[3]{x}-\ln(\cos x)\]\[=\ln \left( \frac{\sqrt[3]{x}(x+1)^2 }{\cos x}\right)\]

OpenStudy (anonymous):

= ln(x+1)^2 + ln x^(1/3) - ln(cosx) = ln [x^(1/3) (x+1)^2] / cos x]